the length of a rectangle is 2 cm greater than twice its width. The area of the rectangle is 40 cm^2. find the dimensions of the rectangle.
Answer by ExplanationCentral(10)

Let the length be l and the width be w.
So we have l = 2w + 2.
The area of the rectangle = length*width = (2w+2)*w = 2w^2 + 2w = 40
Solving this quadratic equation, we get w = 4 cm. This means the length = 2(4) + 2 = 10.
So width = 4 and length = 10.